DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, February 12, 2003 Padelford C-36 3:45 pm

Inverse spectral problems for hyperbolic manifolds

Hiroshi Isozaki (Tokyo Metropolitan University, Japan)

We will discuss some inverse problems for hyperbolic manifolds. As a surprising consequence this leads to new inverse scattering and spectral results in Euclidean space.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.